Supposed a charge is at the origin

Determine the direction angle of the electric field at the point( 0.9, 1.2) measured counterclockwise from the positive x axis if the particle is negatively charged?
1
Expert's answer
2021-09-09T10:51:00-0400


The electric field vector at point ( 0.9, 1.2) is denoted as "\vec{E}" (since the charge is point and negative, it is directed toward the charge). The angle that is required to find is "\theta".

From the right triangle, obtain:


"\beta = 90\degree -\alpha"

In turn:


"\theta = 360\degree -90\degree - \beta = 180+\alpha"

Angle "\alpha" is given as follows:


"\alpha = \arctan\left( \dfrac{1.2}{0.9} \right) \approx 53\degree"

Thus, obtain:


"\theta = 180\degree + 53\degree = 233\degree"

Answer. "233\degree".